ON MAXIMUM MODULUS OF POLYNOMIALS WITH ZEROES OUTSIDE A CIRCLE by V.K.Jain, Kharagpur, India

Abstract. For 0 ≤ r < R ≤ k, the maximum modulus on |z| = r, of a polynomial of degree n, not vanishing in |z| < k, is estimated in terms of its maximum modulus on |z| = R and some of its coefficients. Some other related problems are also considered. Mathematics subject classification: 30A10, 30C10. Keywords: Maximum modulus, zero.